# A Spatio-Temporal Machine Learning Model for Mortgage Credit Risk: Default Probabilities and Loan Portfolios

### v1

@@ Abstract
We introduce a novel machine learning model for credit risk by combining tree-boosting with a latent spatio-temporal Gaussian process model accounting for frailty correlation. This allows for modeling non-linearities and interactions among predictor variables in a flexible data-driven manner and for accounting for spatio-temporal variation that is not explained by observable predictor variables. We also show how estimation and prediction can be done in a computationally efficient manner. In an application to a large U.S. mortgage credit risk data set, we find that both predictive default probabilities for individual loans and predictive loan portfolio loss distributions obtained with our novel approach are more accurate compared to conventional independent linear hazard models and also linear spatio-temporal models. Using interpretability tools for machine learning models, we find that the likely reasons for this outperformance are strong interaction and non-linear effects in the predictor variables and the presence of large spatio-temporal frailty effects.
@@ 1 Introduction
Accurately assessing credit default risk is a challenging and critical task for financial institutions, investors, regulators, and policy makers. Traditionally, linear models such as linear discriminant analysis, logistic regression, or linear discrete hazard models have been used to model default probabilities of individual loans (Altman, 1968; Zmijewski, 1984; Shumway, 2001). Linear hazard models have been extended to account for temporal correlation that cannot be captured by observable covariates (Duffie et al., 2009; Koopman et al., 2011). Such residual correlation is also called frailty correlation. For loans associated with a spatial location, for instance, mortgages and loans to small and medium-sized enterprises (SMEs), there is likely spatial dependence among loans that cannot be fully explained by observable predictor variables, and several linear models have been proposed to account for this spatial frailty correlation (Fernandes and Artes, 2016; Agosto et al., 2019; Calabrese et al., 2019; Babii et al., 2019; Calabrese and Crook, 2020; Medina-Olivares et al., 2022; Calabrese et al., 2024). Recently, linear spatio-temporal models for credit risk have been introduced which model space-time correlation but assume a linear functional form in the predictor variables (Berloco et al., 2023; Medina-Olivares et al., 2023b). However, non-linear machine learning models often achieve a higher prediction accuracy than linear models (Barboza et al., 2017; Zieba et al., 2016; Xia et al., 2017; Sigrist and Hirnschall, 2019; Sigrist and Leuenberger, 2023; Cheraghali and Molnár, 2024). To the best of our knowledge, there exists no prior work that uses state-of-the-art machine learning models and explicitly accounts for spatio-temporal frailty correlation for modeling credit default risk.
@@
In this article, we introduce a novel approach which combines tree-boosting with a latent spatio-temporal Gaussian process. This allows for modeling non-linear and interaction effects of predictor variables as well as for accounting for spatio-temporal frailty correlation among loans which is not accounted for by the observable predictor variables. It is likely that not all relationships are linear, and more realistic non-linear models allow for gaining a better understanding of default mechanisms and for generating more accurate predictions. Furthermore, the space-time Gaussian process allows for generating spatially and temporally varying frailty default risk maps, which can provide valuable insights, while also improving overall prediction accuracy. We apply our proposed model to a large U.S. mortgage data set and compare it to a linear hazard, a linear spatial, and a linear spatio-temporal model regarding the accuracy of predictive default probabilities for individual loans and predictive loan portfolio loss distributions. We first observe that, compared to an independent linear hazard model, incorporating spatial frailty correlation improves the prediction accuracy of default probabilities. Additionally modeling space-time correlations further increases the prediction accuracy compared to a purely spatial linear model. Further, we find that the predictive default probabilities of our proposed spatio-temporal machine learning model are more accurate compared to the default probabilities of a linear spatio-temporal model. Concerning loan portfolios, we analyze the accuracy of predictive loan portfolio loss distributions using the continuous ranked probability score (CRPS) as well as the accuracy of 99% upper quantiles using the corresponding quantile loss. We find that our proposed spatio-temporal frailty machine learning model results in more accurate predictive loan portfolio loss distributions compared to all linear models considered, i.e., a linear hazard model, a linear spatial model, and a linear spatio-temporal model. Interestingly, when considering predictive means and predictive upper quantiles of loan portfolio distributions in the years of the global financial crisis around the year 2009, our tree-boosted spatio-temporal frailty model predicts a higher and thus more realistic loss at the beginning of the crisis, and the loss then reverts faster to lower and more accurate levels after the crisis compared to the other models considered. We also investigate the reasons for the higher prediction accuracy of our proposed spatio-temporal machine learning model, and we find two main explanations. First, there is considerable spatio-temporal variation that is not captured by observable predictor variables. Second, we find that the tree-boosting part of the proposed model accounts for interactions and non-linear effects in the predictor variables that cannot be captured with a linear functional form.
@@
The remainder of this article is organized as follows. In Section 2 we introduce the methodology, and in Section 3, we apply and compare our methodology on a large U.S. mortgage data set. Section 4 concludes.
@@ 2.1 Notation and default probabilities
Our goal is to model default events of $N\in\mathbb{N}^{+}$ loans. For every loan $i$, $i=1,\dots,N$, we assume that we observe predictor variables $X_{ki}\in\mathbb{R}^{p}$ at discrete times $t_{ki}$, $k=0,\dots,n_{i}-1$, and a default time $\tau_{i}\in\{t_{1i},\dots,t_{n_{i}i},\infty\}$, where $0\leq t_{0i}\leq t_{ki}\leq t_{n_{i}i}\leq T$, $k=0,\dots,n_{i}$. I.e., $X_{ki}$ are the predictor variables of loan $i$ observed at time $t_{ki}$, and $\tau_{i}=t_{ki}>t_{0i}$ means that loan $i$ has defaulted in the interval $(t_{k-1i},t_{ki}]$. Further, $t_{0i}$ denotes the time when a loan $i$ enters the set of active loans, $t_{n_{i}i}$ denotes the last observation time for loan $i$, and $n_{i}$ is the total number of temporal observations for loan $i$. The last observation time $t_{n_{i}i}$ can be either the default time $\tau_{i}$, the time of some other form of exit such as reaching the loan maturity date, or the end of the observation period $T$. In addition, we assume that every loan $i$ has associated spatial coordinates $s_{i}\in\mathcal{D}\subset\mathbb{R}^{2}$.
@@ 2.2 Accounting for spatial and spatio-temporal frailty correlation
where $f(\cdot)$ is a link function such as $f(x)=\left(1+e^{-x}\right)^{-1}$, $F(\cdot)$ is a function $F:\mathbb{R}^{p}\mapsto\mathbb{R}$, and the latent variable $b(\cdot,\cdot)$ is a zero-mean Gaussian process (Williams and Rasmussen, 2006) that accounts for spatial or spatio-temporal frailty correlation. Both $b(\cdot,\cdot)$ and $F(\cdot)$ are specified in the following.
For the latent frailty process $b(t,s)$, we consider two cases: a spatial model and a spatio-temporal model. In the spatial model, the Gaussian process $b(t,s)$ varies over space only and is constant over time. It is defined by a spatial covariance function $\text{Cov}(b(t,s),b(t^{\prime},s^{\prime}))=c_{\theta}(s,s^{\prime})$, $s,s^{\prime}\in\mathcal{D}$, which depends on a set of parameters $\theta\in\Theta\subset\mathbb{R}^{q}$. For instance, in the application below, we use a Matérn covariance function
@@
where $\sigma_{1}^{2}$ is a marginal variance parameter, $\rho_{s}$ is a spatial range, or scale, parameter, $\nu$ is a smoothness parameter, $\Gamma(\cdot)$ is the Gamma function, and $K_{\nu}(\cdot)$ is the modified Bessel function of the second kind. For spatio-temporal models, the Gaussian process varies over both space and time, and it is defined by a space-time covariance function $\text{Cov}(b(t,s),b(t^{\prime},s^{\prime}))=c_{\theta}((t,s),(t^{\prime},s^{\prime}))$, $(t,s),(t^{\prime},s^{\prime})\in[0,T]\times\mathcal{D}\subset\mathbb{R}^{3}$. In our application, we use an anisotropic spatio-temporal Matérn covariance function:
@@
where $A=\text{diag}(\rho_{t},\rho_{s},\rho_{s})$ is a diagonal matrix containing a temporal and spatial range parameters $\rho_{t}$ and $\rho_{s}$, respectively.
@@ 2.3 A tree-boosted spatio-temporal Gaussian process model
We assume the model specified in (3) and the subsequent paragraph. In addition, we assume that $F(\cdot)$ is a function in a normed function space $\mathcal{H}$ that is the linear span of a set $\mathcal{S}$ of base learners $f_{j}(\cdot):\mathbb{R}^{p}\rightarrow\mathbb{R}$, which consist of regression trees (Breiman et al., 1984) in this article. Our goal is then to find a joint minimizer for $F(\cdot)\in\mathcal{H}$ and $\theta\in\Theta$ of the functional obtained when plugging $F(\cdot)$ and $\theta$ into the negative log-likelihood:
@@
where $F(X)$ is the row-wise evaluation of $F(\cdot)$ at $X\in\mathbb{R}^{n\times p}$, which is the matrix containing predictor variables for all observations, $n=\sum_{i=1}^{N}n_{i}$, and $\mathcal{L}(F,\theta)$ is given in (6). The minimization of the empirical risk functional in (7) is done iteratively using the latent Gaussian model boosting (LaGaBoost) algorithm given in Algorithm 1, which performs a form of functional gradient descent. In detail, this algorithm iterates between, first, finding a maximum for $\theta$ of $\mathcal{L}(F_{m-1},\theta)$ conditional on the current estimate $F_{m-1}(\cdot)$ and, second, updating the ensemble of trees $F(\cdot)$ using one functional gradient descent step given the current estimates $\theta_{m}$ and $F_{m-1}(\cdot)$. Specifically, the boosting update $f_{m}(\cdot)$ in iteration $m$ is given by the least squares approximation to the vector obtained when evaluating the negative functional gradient of the functional defined in (7) at $(F_{m-1}(\cdot),I_{X_{ki}}(\cdot))$, where $I_{X_{ki}}(\cdot)$ are indicator functions which equal $1$ at $X_{ki}$ and 00 otherwise. Equivalently, $f_{m}(\cdot)$ is the minimizer of a first-order functional Taylor approximation of the functional in (7) with $F(\cdot)=F_{m-1}(\cdot)+f(\cdot)$ around $F_{m-1}(\cdot)$ with an $L^{2}$ penalty on $f(\cdot)$ evaluated at $(X_{ki})$. For more details on the LaGaBoost algorithm, e.g., the calculation of functional gradients; see Sigrist (2022, 2023). In the following, we refer to such a model as “tree-boosted spatio-temporal frailty model” or sometimes shortly as “spatio-temporal LaGaBoost model”.
@@
In order that we can do estimation and prediction in a computationally feasible manner on large data sets in practice, we need to apply some approximations for both the linear Gaussian process and the tree-boosted Gaussian process models. In the following, we describe these approximations. First, the integral in (6) cannot be calculated in closed form, and we approximate it using a Laplace approximation. Laplace approximations are computationally efficient, have asymptotic convergence guarantees, and are accurate for large data sets; see, e.g., Kündig and Sigrist (in press). Furthermore, to ensure that computations with Gaussian processes scale to large data sets, we use Vecchia approximations (Vecchia, 1988; Datta et al., 2016; Katzfuss and Guinness, 2021) for the latent Gaussian processes $b$. In spatial statistics, Vecchia approximations have recently “emerged as a leader among the sea of approximations” (Guinness, 2021) and are often considered as “the most promising class of approximations” (Kang and Katzfuss, 2023). In brief, Vecchia approximations generate an approximate sparse reverse Cholesky factor of the precision matrix of the latent Gaussian process $b$ by using an ordered conditional approximation for the density of the Gaussian process. In doing so, every row of this sparse Cholesky factor contains maximally $m$ non-zero entries corresponding to nearest neighbors for every observation. For prediction, a Vecchia approximation is applied to the joint distribution of a latent Gaussian process at the training and prediction points. Posterior predictive distributions for the latent Gaussian process are then obtained as conditional distributions of this approximated joint distribution. Further, predictive probabilities for the observable response variables are calculated by numerically integrating over the posterior predictive distribution of the latent Gaussian variable using, e.g., adaptive Gauss-Hermite quadrature (Liu and Pierce, 1994). See Sigrist (2023) and Kündig and Sigrist (in press) for more information on Vecchia-Laplace approximations including estimation and prediction. Additionally, for increased computational efficiency, we use the iterative methods of Kündig and Sigrist (in press) for estimation and prediction with Vecchia-Laplace approximations instead of Cholesky decompositions.

### v2

@@ Abstract
We introduce a novel machine learning model for credit risk by combining tree-boosting with a latent spatio-temporal Gaussian process model accounting for frailty correlation. This allows for modeling non-linearities and interactions among predictor variables in a flexible data-driven manner and for accounting for spatio-temporal variation that is not explained by observable predictor variables. We also show how estimation and prediction can be done in a computationally efficient manner. In an application to a large U.S. mortgage credit risk data set, we find that both predictive default probabilities for individual loans and predictive loan portfolio loss distributions obtained with our novel approach are more accurate compared to conventional independent linear hazard models and also linear spatio-temporal models. Using interpretability tools for machine learning models, we find that the likely reasons for this outperformance are strong interaction and non-linear effects in the predictor variables and the presence of spatio-temporal frailty effects.
Keywords: risk analysis, bankruptcy, Gaussian process, space-time frailty correlation, tree-boosting
@@ 1 Introduction
Accurately modeling credit default risk is a challenging and critical task for financial institutions, investors, regulators, and policy makers. Traditionally, linear models such as linear discriminant analysis, logistic regression, and linear discrete hazard models have been widely used to model bankruptcy probabilities in corporate credit risk research (Altman, 1968; Zmijewski, 1984; Shumway, 2001). These linear models are also frequently applied in the retail credit risk literature (Crook et al., 2007; Thomas et al., 2017) for predicting default probabilities of various loan types, including mortgage loans (Agarwal et al., 2012), credit card loans (Bellotti and Crook, 2009b, a, 2013), and personal loans (Banasik et al., 1999; Malik and Thomas, 2010).
For modeling corporate credit risk, linear hazard models have been extended to account for temporal frailty correlation that cannot be captured by observable predictor variables (Duffie et al., 2009; Koopman et al., 2011). Similarly, for loans associated with a specific spatial location, such as mortgages and loans to small and medium-sized enterprises (SMEs), spatial dependence likely exists among loans that cannot be fully explained by observable predictor variables. Specifically, for mortgage credit risk, there are the following potential reasons for the presence of frailty correlation. First, local weather extremes such as wildfires (Issler et al., 2024) and hurricanes (Kousky et al., 2020; Calabrese et al., 2024) have an impact on mortgage credit risk that leads to correlations among defaults. Furthermore, local economic conditions that often cannot be measured using observable predictor variables likely impact default risks. For instance, regions heavily reliant on specific industries are vulnerable to concentrated job losses and income reductions during industry downturns. This can lead to widespread local defaults as borrowers face diminished ability to make payments. In addition, house prices are often locally interconnected, and when local house prices fall, many borrowers in an area can enter negative equity. This can lead to foreclosure contagion which further lowers property values. Harding et al. (2009) show that neighborhood foreclosures have the greatest contagion effect on house prices within a 300-foot radius. This highlights the importance of accounting for local dependencies between properties in immediate neighborhoods.
To account for spatial frailty correlation, several linear models have been proposed in the SME credit risk literature (Fernandes and Artes, 2016; Agosto et al., 2019; Calabrese et al., 2019; Medina-Olivares et al., 2022). Recently, Berloco et al. (2023) introduced a model accounting for spatio-temporal correlation among firms while retaining a linear functional form for the predictor variables. Early approaches to account for spatial dependence in the mortgage credit risk literature include the use of independent and identically distributed frailty variables for metropolitan statistical areas (MSAs) (Kau et al., 2011) and ZIP code concentration variables (Agarwal et al., 2012). Zhu and Pace (2014) demonstrate that spatial probit models offer superior prediction accuracy for mortgage defaults compared to independent probit models. Calabrese and Crook (2020) extended this spatial methodology to linear hazard models. Further research on spatial frailty correlation for mortgage credit risk include Babii et al. (2019); Calabrese et al. (2024), and Medina-Olivares et al. (2023b) who introduce a linear spatio-temporal model to predict mortgage prepayment. Additional prior works on spatial frailty modeling outside the credit risk literature include Li and Ryan (2002), Banerjee and Carlin (2003), Hanson et al. (2012), and Čížek et al. (2017).
Prior research has also investigated the effectiveness of machine learning methods like random forests, tree-boosting, and deep neural networks in modeling corporate credit risk (Barboza et al., 2017; Zieba et al., 2016; Sigrist and Hirnschall, 2019; Sigrist and Leuenberger, 2023; Cheraghali and Molnár, 2024) and retail credit risk (Lessmann et al., 2015; Fitzpatrick and Mues, 2016; Xia et al., 2017; Thomas et al., 2017). These non-linear machine learning methods often outperform traditional linear models in prediction accuracy by capturing complex interactions and non-linear effects in the predictor variables. For instance, Fitzpatrick and Mues (2016) find that tree-boosting is preferred over logistic regression for predicting mortgage defaults. To the best of our knowledge, there exists no prior work that uses state-of-the-art machine learning models and explicitly accounts for spatio-temporal frailty correlation for modeling credit default risk in general and mortgage risk in particular.
@@
In this article, we introduce a novel credit risk modeling approach which combines tree-boosting with a latent spatio-temporal Gaussian process. This allows for modeling non-linear and interaction effects of predictor variables as well as for accounting for spatio-temporal frailty correlation among loans which is not accounted for by the observable predictor variables. It is likely that not all relationships are linear, and more realistic non-linear models allow for gaining a better understanding of default mechanisms and for generating more accurate predictions. Furthermore, a space-time Gaussian process integrated in the model allows for accounting for frailty correlation and for generating spatially and temporally varying frailty default risk maps. Besides providing valuable insights on its own, this can improve the prediction accuracy of default probabilities and loan portfolio distributions. To the best of our knowledge, this is the first work combining space-time Gaussian processes with tree-boosting, also outside of the credit risk literature.
We apply our proposed model to a large U.S. mortgage data set and compare it to a linear hazard, a linear spatial, and a linear spatio-temporal model. We first observe that incorporating spatial and spatio-temporal frailty correlation in a linear model improves the prediction accuracy of default probabilities compared to an independent linear hazard model. Our newly proposed methodology, which extends spatial and spatio-temporal linear models using tree-boosting, leads to further improvements in prediction accuracy of default probabilities. Although the differences in prediction accuracy for individual default probabilities are relatively small, they are nonetheless important, as even small differences are relevant for large portfolios of mortgages. For example, when considering loan portfolios consisting of all active mortgages in our data set, we find that our proposed spatial and spatio-temporal frailty machine learning models yield considerably more accurate predictive 99% upper quantiles in terms of the quantile loss compared to all linear models, i.e., a linear hazard model without frailty correlation, a linear spatial model, and a linear spatio-temporal frailty model. In particular, the proposed tree-boosted frailty models predict more realistic losses at the beginning and end of the global financial crisis around the year 2008 with, e.g., mean losses of frailty machine learning models being almost US \$1 billion more accurate compared to linear models. Using interpretability tools for machine learning models, we find that there are strong interactions and non-linear effects in the predictor variables which cannot be captured with a linear functional form. In addition, we find that there are sizable spatio-temporal frailty effects.
@@
The remainder of this article is organized as follows. In Section 2, we introduce the methodology, and in Section 3, we apply and compare our methods on a large U.S. mortgage data set. Section 4 concludes.
@@ 2.1 Notation and default probabilities
Our goal is to model default events of $N\in\mathbb{N}^{+}$ loans. For every loan $i$, $i=1,\dots,N$, we observe predictor variables $X_{ki}\in\mathbb{R}^{p}$ at discrete times $t_{ki}$, $k=0,\dots,n_{i}-1$, and a default time $\tau_{i}\in\{t_{1i},\dots,t_{n_{i}i},\infty\}$, where $0\leq t_{0i}\leq t_{ki}\leq t_{n_{i}i}\leq T$, $k=0,\dots,n_{i}$. I.e., $X_{ki}$ are the predictor variables of loan $i$ observed at time $t_{ki}$, and $\tau_{i}=t_{ki}>t_{0i}$ means that loan $i$ has defaulted in the interval $(t_{k-1i},t_{ki}]$. Further, $t_{0i}$ denotes the time when a loan $i$ enters the set of active loans, $t_{n_{i}i}$ denotes the last observation time for loan $i$, and $n_{i}$ is the total number of temporal observations for loan $i$. The last observation time $t_{n_{i}i}$ can be either the default time $\tau_{i}$, the time of some other form of exit such as reaching the loan maturity date, or the end of the observation period $T$. In addition, we assume that every loan $i$ has associated spatial coordinates $s_{i}\in\mathcal{D}\subset\mathbb{R}^{2}$.
@@ 2.2 Accounting for spatial and spatio-temporal frailty correlation
where $f(\cdot)$ is a link function such as $f(x)=\left(1+e^{-x}\right)^{-1}$, $F(\cdot)$ is a function $F:\mathbb{R}^{p}\mapsto\mathbb{R}$, and the latent variable $b(\cdot,\cdot)$ is a zero-mean Gaussian process (Rasmussen and Williams, 2006) that accounts for spatial or spatio-temporal frailty correlation. Both $b(\cdot,\cdot)$ and $F(\cdot)$ are specified in the following.
For the latent frailty process $b(t,s)$, we consider two cases: a spatial model and a spatio-temporal model. In the spatial model, the Gaussian process $b(t,s)$ varies over space only and is constant over time. It is defined by a spatial covariance function $\text{Cov}(b(t,s),b(t^{\prime},s^{\prime}))=c_{\theta}(s,s^{\prime})$, $s,s^{\prime}\in\mathcal{D}$, which depends on a set of parameters $\theta\in\Theta\subset\mathbb{R}^{q}$. For more information on covariance functions in general, see Rasmussen and Williams (2006, Chapter 4), and for spatial covariance functions in particular, see Cressie (2015, Section 2.3.1). In the application below, we use a Matérn covariance function:
@@
where $\sigma_{1}^{2}$ is a marginal variance parameter, $\rho_{s}$ is a spatial range, or scale, parameter, $\nu$ is a smoothness parameter, $\Gamma(\cdot)$ is the Gamma function, and $K_{\nu}(\cdot)$ is the modified Bessel function of the second kind. For spatio-temporal models, the Gaussian process varies over both space and time, and it is defined by a space-time covariance function $\text{Cov}(b(t,s),b(t^{\prime},s^{\prime}))=c_{\theta}((t,s),(t^{\prime},s^{\prime}))$, $(t,s),(t^{\prime},s^{\prime})\in[0,T]\times\mathcal{D}\subset\mathbb{R}^{3}$. For more information on space-time covariance functions, we refer to Cressie and Wikle (2011, Section 6.1). In our application, we use an anisotropic spatio-temporal Matérn covariance function:
@@
where $A=\text{diag}(\rho_{t},\rho_{s},\rho_{s})$ is a diagonal matrix containing a temporal and spatial range parameters $\rho_{t}$ and $\rho_{s}$, respectively. We mainly choose this space-time covariance function since it is a parsimonious model that allows for computationally efficient inference when using a Vecchia approximation and a correlation-based neighbor selection approach (Kang and Katzfuss, 2023), see below in Section 3.3 for more details.
@@ 2.3 A tree-boosted spatio-temporal Gaussian process model
We assume the model specified in (3) in the subsequent paragraph. In addition, we assume that $F(\cdot)$ is a function in a normed function space $\mathcal{H}$ that is the linear span of a set $\mathcal{S}$ of base learners $f_{j}(\cdot):\mathbb{R}^{p}\rightarrow\mathbb{R}$, which consist of regression trees (Breiman et al., 1984) in this article. Our goal is then to find a joint minimizer for $F(\cdot)\in\mathcal{H}$ and $\theta\in\Theta$ of the functional obtained when plugging $F(\cdot)$ and $\theta$ into the negative log-likelihood:
@@
where $F(X)$ is the row-wise evaluation of $F(\cdot)$ at $X\in\mathbb{R}^{n\times p}$, which is the matrix containing predictor variables for all observations, $n=\sum_{i=1}^{N}n_{i}$, and $\mathcal{L}(F,\theta)$ is given in (6). The minimization of the empirical risk functional in (7) is done iteratively using the latent Gaussian model boosting (LaGaBoost) algorithm given in Algorithm 1, which performs a form of functional gradient descent. In detail, this algorithm iterates between, first, finding a maximum for $\theta$ of $\mathcal{L}(F_{m-1},\theta)$ conditional on the current estimate $F_{m-1}(\cdot)$ and, second, updating the ensemble of trees $F(\cdot)$ using one functional gradient descent step given the current estimates $\theta_{m}$ and $F_{m-1}(\cdot)$. Specifically, the boosting update $f_{m}(\cdot)$ in iteration $m$ is given by the least squares approximation to the vector obtained when evaluating the negative functional gradient of the functional defined in (7) at $(F_{m-1}(\cdot),I_{X_{ki}}(\cdot))$, where $I_{X_{ki}}(\cdot)$ are indicator functions which equal $1$ at $X_{ki}$ and 00 otherwise. Equivalently, $f_{m}(\cdot)$ is the minimizer of a first-order functional Taylor approximation of the functional in (7) with $F(\cdot)=F_{m-1}(\cdot)+f(\cdot)$ around $F_{m-1}(\cdot)$ with an $L^{2}$ penalty on $f(\cdot)$ evaluated at $(X_{ki})$. For more details on the LaGaBoost algorithm, e.g., the calculation of functional gradients; see Sigrist (2022, 2023).
@@
In order that we can do estimation and prediction in a computationally feasible manner on large data sets in practice, we need to apply some approximations for both the linear Gaussian process and the tree-boosted Gaussian process models. In the following, we describe these approximations. First, the integral in (6) cannot be calculated in closed form, and we approximate it using a Laplace approximation. Laplace approximations are computationally efficient, have asymptotic convergence guarantees, and are accurate for large data sets; see, e.g., Kündig and Sigrist (2025). Furthermore, to ensure that computations with Gaussian processes scale to large data sets, we use Vecchia approximations (Vecchia, 1988; Datta et al., 2016; Katzfuss and Guinness, 2021) for the latent Gaussian process $b$. In spatial statistics, Vecchia approximations have recently “emerged as a leader among the sea of approximations” (Guinness, 2021) and are often considered as “the most promising class of approximations” (Kang and Katzfuss, 2023; Rambelli and Sigrist, 2025). In brief, Vecchia approximations calculate an approximate sparse reverse Cholesky factor of the precision matrix of the latent Gaussian process $b$ by using an ordered conditional approximation for the density of the Gaussian process. In doing so, every row of this sparse Cholesky factor contains maximally $\tilde{m}$ non-zero entries. For prediction, a Vecchia approximation is applied to the joint distribution of a latent Gaussian process at the training and prediction points. Posterior predictive distributions for the latent Gaussian process are then obtained as conditional distributions of this approximated joint distribution. Further, predictive probabilities for the observable response variables are calculated by numerically integrating over the posterior predictive distribution of the latent Gaussian variable using, e.g., adaptive Gauss-Hermite quadrature (Liu and Pierce, 1994). We describe the construction of Vecchia approximations for the latent Gaussian process in more detail in Appendix A.1. For more information on Vecchia-Laplace approximations, including estimation and prediction, see Sigrist (2023) and Kündig and Sigrist (2025). Additionally, for increased computational efficiency, we use the iterative methods of Kündig and Sigrist (2025) for estimation and prediction with Vecchia-Laplace approximations instead of Cholesky decompositions.